OpenStudy (anonymous):

exponential equation: 5^(1-2x) = 1/5 I've started out with (1-2x)log(5)=log1/5 Is this the right path?

OpenStudy (campbell_st):

1/5 - 5^(-1) so equating powers 1 - 2x = -1 -2x = -2 x = 1

OpenStudy (kinggeorge):

\[5^{1-2x}={1 \over 5}=5^{-1}\]Take log base 5 of both sides.\[\log_5{5^{1-2x}}=\log_5{5^{-1}}\]So you get\[1-2x=-1\]Thus, \(x=1\).

OpenStudy (anonymous):

oooooh ok, I totlaly forgot about that rule. It's coming back now. I was getting laws mixed up. Too many to remember! I know exactly what you're talking about now.

OpenStudy (anonymous):

Ok Yes you are doing it right!!! From where you left off. (1-2x)log(5)=log(1/5) (1-2x)=log(1/5)/log(5) -----Divide the log(5) (1-2x)= -1 -----log(1/5)/log(5) is just -1 1-2x-1= -1-1 -2x= -2 x= 1

OpenStudy (anonymous):

@Romero awesome, thank you. this helps tremendouly. Logs have been my enemy this semester!!!